OpenStudy (anonymous):

What is the radius for the circle given by the equation x^2+(y-3)^2=21? Round your answer to the nearest thousandth.

OpenStudy (phi):

do you know the equation of a circle ?

OpenStudy (anonymous):

No, do you mind helping me out with that?

OpenStudy (phi):

see http://www.mathsisfun.com/algebra/circle-equations.html and scroll down to see the equation

OpenStudy (phi):

can you write down the equation of a circle ?

OpenStudy (anonymous):

(x-a)^2+(y-b)^2-r^2. is that correct?

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2; correct one a,b are the coordinates of its center and r is radius so just compare those two equations

OpenStudy (anonymous):

How do I find the radius and coordinates though?

OpenStudy (anonymous):

comparison of thq eqn in given ques and the the eqn above

OpenStudy (anonymous):

Sooo 0 and 3 would be coordinates and 21 is the radius?

OpenStudy (anonymous):

noo...rt21 is the radius as 21=r^2

OpenStudy (anonymous):

okay, so square root of 21?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

But it says "Round your answer to the nearest thousandth." how am I supposed to do that?

OpenStudy (anonymous):

use calculator or ancient way to find root

OpenStudy (anonymous):

I got 4.58257569... so would the answer be 4.59?

OpenStudy (phi):

the thousandth digit is the third digit to the right of the decimal point look at the 4th digit to the right. if smaller than 5, just keep the 3 digits as is if the the 4th digit is 5 or more, add 1 to the 3rd digit.

OpenStudy (anonymous):

thousands got 3 digits so 4.523

OpenStudy (anonymous):

wouldn't it be 4.583?

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you! :)